# Learning Beyond Experience: Generalizing to Unseen State Space with Reservoir Computing

### v1

@@ V Discussion
These findings suggest that RCs can construct a flexible internal representation of system dynamics that extends beyond the training data. In doing so, they can capture global structures such as basins of attraction and make reliable forecasts in regions of state space that were never directly observed. This kind of generalization – achieved from sparse, disjoint training data and without structural assumptions – positions RCs as a practical tool for modeling complex systems in settings where prior knowledge is limited or unavailable.
@@ Acknowledgements.
The code that supports the findings of this study are available at the following repository:
@@ Appendix A Experimental Setup
We obtain trajectories of the Duffing system by integrating Eq. 13 using a fourth order Runge-Kutta integrator. We generate trajectories of the magnetic pendulum system by integrating Eq. 16 using using the scipy integrate.solve_ivp implementation of the DOP853 eigth order Runge-Kutta integration scheme. For the Duffing system, the integration time step is fixed (${\Delta t=0.01}$) and for the magnetic pendulum it is adaptive. In both cases, the time step between samples in the RCs’ training and test signals are fixed: ${\Delta t=0.01}$ for the Duffing system and ${\Delta t=0.02}$ for the magnetic pendulum, as in Table 1. Because we intend for the RC to learn from and predict the transient dynamics of each system, we do not discard any portion of the integrated trajectories before forming training or test signals. (We do however, discard a short transient response of the reservoir’s internal state at the start of each training signal, as described in Section III.)
Duffing Magnetic
Time Step $\Delta t$ $0.01$ $0.02$
Training Transient $N_{trans}$ 5 25
Training Signal Length $N_{i}\ \forall\ i$ 500 500
Forecast Horizon $N_{f}$ 2000 2000
Distance Threshold $\varepsilon_{c}$ 0.5 0.25
@@
The parameters we use to construct and train RCs for our experiments with the Duffing and magnetic pendulum systems are provided in Table 2. The other parameters defining our experiments are in Table 1. For simplicity, we use training trajectories that are of equal length, ${N_{i}=500\ \forall\ i=1,...,N_{train}}$, in all of our experiments. Our multiple-trajectory training scheme (Section III) does not, however, require that all training signals are the same length. When all training trajectories must be from the same basin of attraction, we first generate a trajectory of length $4000\Delta t$ from each sample initial condition and check whether the trajectory converges to the corresponding desired attractor. If a trajectory converges to the right attractor, we include its first ${N_{i}=500}$ data points in the RC’s training data. We repeat this process with new sample initial conditions until we obtain the desired number of training trajectories, $N_{train}$, from the chosen basin. In all of our experiments, we forecast from the end of each provided test signal, i.e., ${t=(N_{test}-1)\Delta t}$, to time ${t=N_{f}\Delta t=2000\Delta t}$.
As shown in Table 2, the reservoir hyperparameters that we use for our experiments with the Duffing and magnetic pendulum systems are identical except for the input strength range, $\sigma$, regularization strength, $\alpha$, and noise amplitude, $\eta$. We chose these hyperparameters by coarse hand-tuning to allow for good, but not necessarily optimal performance. We chose the other hyperparemeters to have values that typically allow for reasonably accurate forecasting with reservoir computers, and performed no experiment-specific tuning of these values. While more robust hyperparameter tuning may improve performance overall, our priority is to demonstrate that reservoir computers can generalize to unexplored regions of state space without system-specific structural constraints, rather than to obtain highly optimized forecasts.
Duffing Magnetic
Reservoir Size $N_{r}$ 75 2500
Mean In-degree $\langle d\rangle$ 3 3
Input Strength Range $\sigma$ 1.0 5.0
Spectral Radius $\rho$ 0.4 0.4
Bias Strength Range $\psi$ 0.5 0.5
Leakage Rate $\lambda$ 1.0 1.0
Tikhonov Regularization $\alpha$ $10^{-8}$ $10^{-6}$
Training Noise Amplitude $\eta$ $10^{-5}$ $10^{-3}$
@@
Figure 7: RCs can reliably generalize to an unseen basin of the Duffing system, unless the training data are very restricted. We train a reservoir computer (RC) of ${N_{r}=75}$ nodes to forecast fully-observed states of the Duffing system. We draw ${N_{train}=10}$ training signals (top row) from the basin of attraction ${B(\bm{A}^{-})}$ for the fixed point ${\bm{A}^{-}}$, and make predictions from $36$ test signals of length ${N_{test}=10}$ points, whose initial conditions are drawn randomly from ${-10\leq x_{0}\leq 10}$ and ${-10\leq y_{0}\leq 10}$ and from both basins of attraction, ${B(\bm{A}^{\pm})}$. We restrict the training initial conditions to the ranges ${-\Delta_{0}^{train}\leq x_{0}\leq\Delta_{0}^{train}}$ and ${-\Delta_{0}^{train}\leq y_{0}\leq\Delta_{0}^{train}}$. In (a) to (d) we decrease the half-width of the range of the training initial conditions, from ${\Delta_{0}^{train}=10}$ (a) to ${\Delta_{0}^{train}=4}$ (d), while holding the half-width of the range of the test initial conditions, ${\Delta_{0}^{test}=10}$, fixed. As ${\Delta_{0}^{train}}$ decreases, the RC’s performance away from the fixed points diminishes. When ${\Delta_{0}^{train}=4}$, the RC predicts the existence of a spurious attractor, which does not coincide with the position of the true unseen attractor, ${\bm{A}^{+}\approx(3.16,0)}$.
@@
Figure 8: RC Basin Predictions and Comparison with Baseline Across Test Signal Lengths. Row 1: Reservoir computer (RC) basin predictions over a ${300\times 300}$ grid of initial conditions, i.e., the state at the start of the test signal. Each point is colored by the attractor to which the RC predicts convergence (pink, blue, yellow). Green points indicate that the RC predicts convergence to a spurious attractor. Fractions correctly classified: $f_{c}\approx 0.54$ ($N_{test}=25$), $f_{c}\approx 0.57$ ($N_{test}=50$), $f_{c}\approx 0.67$ ($N_{test}=100$). Row 2: Baseline predictions over the same grid of initial conditions, where each point is assigned to the nearest fixed point based on its state at the end of the test signal (i.e., the start of the prediction interval). Fractions correct: ${f_{c}\approx 0.45}$ ($N_{test}=1$, not plotted), ${f_{c}\approx 0.45}$ ($N_{test}=25$), ${f_{c}\approx 0.45}$ (${N_{test}=50}$), ${f_{c}\approx 0.47}$ (${N_{test}=100}$). Rows 1 and 2 show full-color basin maps spanning the entire initial condition grid. Rows 3–5: Predictions and ground truth at the locations of the measured state at the end of each test signal. These panels contain only discrete points (colored dots) at those final states, with white backgrounds elsewhere. Row 3: Baseline predictions for the final test signal states. Row 4: RC predictions for the final test signal states. Row 5: Ground-truth basin assignments at the final states of the test signals. When visualizing and comparing the RC predictions (row 4) vs ground truth (row 5) at the final states of the test signals, we see a similar trend as we saw when visualizing the basin maps from the locations of the initial conditions (Fig. 5 and row 1): the RC predictions and the ground truth have qualitatively similar basin structures. These plots, along with the calculated fractions correct, also indicate that the baseline method – assigning each point to the nearest basin at the end of the test signal – shows only modest improvement across the test signal lengths considered.
@@
Figure 9: In many scenarios, an RC generalizing to unseen basins of attraction of the magnetic pendulum performs comparably to one trained on data from all three of its basins. We train an RC of ${N_{r}=2500}$ nodes to forecast partially-observed states of the magnetic pendulum system (the position components $x$ and $y$ only) and plot the fraction of initial conditions for which the RC predicts the correct basin of attraction as we vary the number of training initial conditions, $N_{train}$, and the half-widths of the ranges of the training and test initial conditions, ${\Delta_{0}^{train}}$ and ${\Delta_{0}^{test}}$. In (a) and (c), we draw all training signals from the basin of attraction of the magnet at ${(\tilde{x}_{1},\tilde{y}_{1})=(1/\sqrt{3},0)}$, and in (c) and (d), we draw training signals from all basins of attraction. In all cases, we randomly sample the initial conditions of the training trajectories, but the initial conditions of the test signals (of length ${N_{test}=100}$), which belong to both basins of attraction, form a ${50\times 50}$ uniform grid. For all grid points, we plot the mean fraction correct calculated over five random realizations of the RC’s internal structure and of the initial conditions of the training signals. In (a) and (b), we fix the number of training initial conditions, ${N_{train}=100}$, and in (c) and (d), we fix the half-width of the test initial condition range, ${\Delta_{0}^{test}=1.5}$. Vertical lines mark the distance of the fixed points from the origin and diagonal lines mark ${\Delta_{0}^{test}=\Delta_{0}^{train}}$. Red crosses indicate the values of $N_{train}$, $\Delta_{0}^{train}$, and $\Delta_{0}^{test}$ that we use in Figs. 5, 6 and 8.

### v2

@@ V Discussion
These findings suggest that RCs can construct a generalizable internal representation of system dynamics that extends beyond the training data. In doing so, they can capture global structures such as basins of attraction and make reliable forecasts in regions of state space that were never directly observed. This kind of generalization – achieved from sparse, disjoint training data and without structural assumptions – positions RCs as a practical tool for modeling complex systems in settings where prior knowledge is unavailable and data are limited.
@@
Supplementary Material
See the supplementary material for Figs.˜S1, S2, S3, S5 and S4, which provide further information on the sensitivity of RC generalization to dynamical/process noise, test signal length, and training data diversity, as well as illustrations of predicted near-attractor behavior and spurious attractors.
@@ Acknowledgements.
Author Contributions
D. A. Norton: Conceptualization (equal); Formal analysis (lead); Funding acquisition (equal); Investigation (lead); Methodology (equal); Software (lead); Validation (lead); Visualization (lead); Writing – original draft (lead). Y. Zhang: Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Visualization (supporting); Writing – review & editing (equal). M. Girvan: Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Project administration (lead); Resources (lead); Supervision (lead); Visualization (supporting); Writing – review & editing (equal).
@@
The code that supports the findings of this study is available at the following repository:
@@ Appendix A Experimental Setup
We obtain trajectories of the Duffing, multi-well, and multistable Lorenz-like systems by integrating Eqs.˜13, 16 and 19 using a fourth order Runge-Kutta integrator. We generate trajectories of the magnetic pendulum system by integrating Eq.˜17 using using the scipy integrate.solve_ivp implementation of the DOP853 eigth order Runge-Kutta integration scheme. For the magnetic pendulum, the integration time step is adaptive and for the other systems, it is fixed ($\Delta t=0.01$}). In all cases, the time step between samples in the RCs’ training and test signals is fixed: $\Delta t=0.01$} for the Duffing and multi-well systems and $\Delta t=0.02$} for the magnetic pendulum and multistable Lorenz-like systems, as in Table˜2. Because we intend for the RC to learn from and predict the transient dynamics of each system, we do not discard any portion of the integrated trajectories before forming training or test signals. (We do however, discard a short transient response of the reservoir’s internal state at the start of each training signal, as described in Section˜III.)
@@
Duffing Wells Magnetic Lorenz
RC Time Step $\Delta t$ $0.01$ $0.01$ $0.02$ $0.02$
Training Transient $N_{trans}$ 5 5 25 5
Training Signal Length $N_{i}\ \forall\ i$ 500 500 500 5000
Forecast Horizon $N_{f}$ 2000 2000 2000 5000
Convergence Threshold $\varepsilon_{c}$ 0.5 0.25 1.0
The parameters we use to construct and train RCs for our experiments are provided in Table˜3. The other parameters defining our experiments are in Table˜2. For simplicity, we use training trajectories that are of equal length, $N_{i}=500\ \forall\ i=1,...,N_{train}$}, in all of our experiments (except for the multistable Lorenz-like system, which we address separately in the next paragraph). Our multiple-trajectory training scheme (Section˜III) does not, however, require that all training signals are the same length. When all training trajectories must be from the same basin of attraction, we first generate a trajectory of length $4000\Delta t$ from each sample initial condition and check whether the trajectory converges to the corresponding desired attractor. If a trajectory converges to the right attractor, we include its first $N_{i}=500$} data points in the RC’s training data. We repeat this process with new sample initial conditions until we obtain the desired number of training trajectories, $N_{train}$, from the chosen basin. In all of our experiments (except the Lorenz-like system), we forecast from the end of each provided test signal, i.e., $t=(N_{test}-1)\Delta t$}, to time $t=N_{f}\Delta t=2000\Delta t$}.
@@
Duffing Wells Magnetic Lorenz
Reservoir Size $N_{r}$ 200 200 2500 500
Mean In-degree $\langle d\rangle$ 10 10 10 10
Input Strength Range $\sigma$ 1.0 1.0  [^1] We use $\sigma=1.0$} when we standardize inputs to the reservoir and $\sigma=0.25$} when we do not standardize the inputs. 5.0 0.5
Spectral Radius $\rho$ 0.4 0.4 0.4 0.4
Bias Strength Range $\psi$ 0.5 0.5 0.5 0.5
Leakage Rate $\lambda$ 1.0 1.0 1.0 1.0
Tikhonov Regularization $\alpha$ $10^{-12}$ $^{-12} 1010$$^{-10}$ $10^{-10}$
Training Noise Amplitude $\eta$ $10^{-5}$ $10^{-5}$ $10^{-3}$ $10^{-3}$
For our experiments with the multistable Lorenz-like system, we use only one training trajectory, which is of length $N_{1}=5000$} to allow for adequate sampling of the on-attractor chaotic dynamics. Similarly, for each test signal, we forecast from $t=(N_{test}-1)\Delta t$}, to time $t=N_{f}\Delta t=5000\Delta t$}. To assess whether a prediction converges to one of the system’s true attractors, we measure the Kullback-Liebler (KL) divergence of the empirical distribution of Lorenz states over the final $500$ time steps of the predicted trajectory relative to that of a reference trajectory on the attractor, as described in Section˜IV.4. We measure the KL divergences using only the final $500$ time steps of the predicted trajectory to ensure that (1) the transient approach to the attractor does not obscure the on-attractor behavior and (2) the RC’s prediction is stable and has remained on the attractor for a prolonged period (the previous $4500-N_{test}$} predicted data points). To estimate the KL divergence, we use the dystsGilpin (2021, 2023) metrics.estimate_kl_divergence routine with its default arguments: $\text{n\_samples}=1000$}, $\text{sigma\_scale}=1$}, and $\text{eps}=10^{-10}$}.
As shown in Table˜3, the reservoir hyperparameters that we use for our experiments with the Duffing, multi-well, magnetic pendulum, and Lorenz-like systems are identical except for the input strength range, $\sigma$, regularization strength, $\alpha$, and noise amplitude, $\eta$. We chose these hyperparameters by coarse hand-tuning to allow for good, but not necessarily optimal performance. We chose the other hyperparameters to have values that typically allow for reasonably accurate forecasting with reservoir computers, and performed no experiment-specific tuning of these values. While more robust hyperparameter tuning may improve performance overall, our priority is to demonstrate that reservoir computers can generalize to unexplored regions of state space without system-specific structural constraints, rather than to obtain highly optimized forecasts.
@@
Figure S1: RCs can reliably generalize to an unseen basin of the Duffing system, unless the training data are very restricted. We train a reservoir computer (RC) of $N_{r}=200$} nodes to forecast fully-observed states of the Duffing system. We draw $N_{train}=10$} training signals (top row) from the basin of attraction $B(\bm{A}^{-})$} for the fixed point $\bm{A}^{-}$}, and make predictions from $36$ test signals of length $N_{test}=10$} points, whose initial conditions are drawn randomly from $10\leq x_{0}\leq 10$} and $10\leq y_{0}\leq 10$} and from both basins of attraction, $B(\bm{A}^{\pm})$}. We restrict the training initial conditions to the ranges $\Delta_{0}^{train}\leq x_{0}\leq\Delta_{0}^{train}$} and $\Delta_{0}^{train}\leq y_{0}\leq\Delta_{0}^{train}$}. In (a) to (d) we decrease the half-width of the range of the training initial conditions, from $\Delta_{0}^{train}=10$} (a) to $\Delta_{0}^{train}=4$} (d), while holding the half-width of the range of the test initial conditions, $\Delta_{0}^{test}=10$}, fixed. As $\Delta_{0}^{train}$} decreases, the RC’s performance away from the fixed points diminishes. When $\Delta_{0}^{train}=4$}, the RC predicts the existence of a spurious attractor, which does not coincide with the position of the true unseen attractor, $\bm{A}^{+}\approx(3.16,0)$}.
Figure S2: RCs can reliably generalize to an unseen basin of the Duffing system, even when moderate stochastic forcing (dynamical/process noise) influences the training trajectories. We train an RC of $N_{r}=200$} nodes on $N_{train}=10$} partially-observed trajectories ($x$ only) of the unforced Duffing system from the basin of attraction $B(\bm{A}^{-})$ (blue). We integrate dynamical/process noise into the training trajectories by including white noise in the Duffing system’s equations of motion: $\dot{x}=y+\mathcal{N}(0,\eta_{p})$}, $\dot{y}=ay-bx-cx^{3}+\mathcal{N}(0,\eta_{p})$}, where $\mathcal{N}(0,\eta_{p})$} draws a random variable from a Gaussian distribution with mean zero and standard deviation $\eta_{p}$. The noise is independent for each component of the system and at each integration time step. We then make predictions from short noiseless test signals ($N_{test}=10$} observations each) from both basins of attraction. (a) to (d): The RC-predicted basins for different values of $\eta_{p}$, from $\eta_{p}=0.1$} (a) to $\eta_{p}=2.0$} (d). The blue/pink initial conditions are those that the RC correctly predicts belong to the seen/unseen basin. The initial conditions of test signals for which the RC-predicted trajectory converges to the incorrect attractor are yellow. Crosses mark the true fixed point attractors and dots mark the training initial conditions. We use $y_{0}$ for plotting purposes only; the RC has access to and then predicts only the $x$-component of the Duffing system.
@@
Figure S3: RC Basin Predictions and Comparison with Baseline Across Test Signal Lengths. Row 1: Reservoir computer (RC) basin predictions over a $300\times 300$} grid of initial conditions, i.e., the state at the start of the test signal. Each point is colored by the attractor to which the RC predicts convergence (pink, blue, yellow). Green points indicate that the RC predicts convergence to a spurious attractor. Fractions correctly classified: $f_{c}\approx 0.52$} ($N_{test}=25$}), $f_{c}\approx 0.57$} ($N_{test}=50$}), $f_{c}\approx 0.66$} ($N_{test}=100$}). Row 2: Baseline predictions over the same grid of initial conditions, where each point is assigned to the nearest fixed point based on its state at the end of the test signal (i.e., the start of the prediction interval). Fractions correct: $f_{c}\approx 0.45$} ($N_{test}=1$}, not plotted), $f_{c}\approx 0.45$} ($N_{test}=25$}), $f_{c}\approx 0.45$} ($N_{test}=50$}), $f_{c}\approx 0.47$} ($N_{test}=100$}). Rows 1 and 2 show full-color basin maps spanning the entire initial condition grid. Rows 3–5: Predictions and ground truth at the locations of the measured state at the end of each test signal. These panels contain only discrete points (colored dots) at those final states, with white backgrounds elsewhere. Row 3: Baseline predictions for the final test signal states. Row 4: RC predictions for the final test signal states. Row 5: Ground-truth basin assignments at the final states of the test signals. When visualizing and comparing the RC predictions (row 4) vs ground truth (row 5) at the final states of the test signals, we see a similar trend as we saw when visualizing the basin maps from the locations of the initial conditions (Fig.˜6 and row 1): the RC predictions and the ground truth have qualitatively similar basin structures. These plots, along with the calculated fractions correct, also indicate that the baseline method – assigning each point to the nearest basin at the end of the test signal – shows only modest improvement across the test signal lengths considered.
@@
Figure S4: In many scenarios, an RC generalizing to unseen basins of attraction of the magnetic pendulum performs comparably to one trained on data from all three of its basins. We train an RC of $N_{r}=2500$} nodes to forecast partially-observed states of the magnetic pendulum system (the position components $x$ and $y$ only) and plot the fraction of initial conditions for which the RC predicts the correct basin of attraction as we vary the number of training initial conditions, $N_{train}$, and the half-widths of the ranges of the training and test initial conditions, $\Delta_{0}^{train}$} and $\Delta_{0}^{test}$}. In (a) and (c), we draw all training signals from the basin of attraction of the magnet at $(\tilde{x}_{1},\tilde{y}_{1})=(1/\sqrt{3},0)$}, and in (c) and (d), we draw training signals from all basins of attraction. In all cases, we randomly sample the initial conditions of the training trajectories, but the initial conditions of the test signals (of length $N_{test}=100$}), which belong to both basins of attraction, form a $50\times 50$} uniform grid. For all grid points, we plot the mean fraction correct calculated over ten random realizations of the RC’s internal structure and of the initial conditions of the training signals. In (a) and (b), we fix the number of training initial conditions, $N_{train}=100$}, and in (c) and (d), we fix the half-width of the test initial condition range, $\Delta_{0}^{test}=1.5$}. Vertical lines mark the distance of the fixed points from the origin and diagonal lines mark $\Delta_{0}^{test}=\Delta_{0}^{train}$}. Red crosses indicate the values of $N_{train}$, $\Delta_{0}^{train}$, and $\Delta_{0}^{test}$ that we use in Figs.˜6, S5 and S3.
Figure S5: RCs may learn some attractors more accurately than others, even as they generalize to unseen basins. (a) to (c) We measure the prediction error (Euclidean distance between the true and predicted trajectories) over time for sample predictions from short test signals ($N_{test}=100$} data points) in the seen pink basin (a) and in the unseen blue (b) and yellow (c) basins. (d) Distributions of the maximum errors over the final $25$ predicted points, calculated over $900$ predictions from test signals with randomly-distributed initial conditions ($310$ pink, $304$ blue, $286$ yellow). We color lines (a to c) according to their predicted basins and histograms (d) according to their true basins. In all panels, dashed lines indicate the threshold distance used to establish convergence, $\varepsilon_{c}=0.25$. We use the same RC and $N_{train}=100$ training trajectories from the pink basin of attraction as in Fig.˜6. The RC learns the dynamics near the seen pink attractor slightly more accurately than near the unseen blue and yellow attractors.